# Universal Dense-Matter Trace Anomaly Inferred from Collective Flow in Heavy-Ion Collisions and Global Properties of Neutron Stars

### v1

@@ Abstract
The trace anomaly of dense matter, $\Delta\equiv 1/3-P/\varepsilon$, defined in terms of the ratio of pressure $P$ to energy density $\varepsilon$, quantifies deviations from conformal symmetry and plays a central role in both the hydrodynamic response and gravitational equilibrium. While $\Delta(\varepsilon)$ has recently been inferred from neutron star observations, we report the first Bayesian extraction of the trace anomaly from collective flow observables in intermediate-energy heavy-ion collisions. By employing transport-model simulations that explicitly decouple the cold-matter mean-field potential from thermal effects, we directly constrain the cold dense-matter equation of state (EOS). Remarkably, the trace anomaly inferred from laboratory flow data agrees quantitatively, within $68\%$ credible intervals, with independent astrophysical posterior bands. This nontrivial agreement demonstrates that heavy-ion collisions and neutron star observations probe the same universal macroscopic properties of dense matter, establishing the trace anomaly as a composition-insensitive descriptor of dense matter across widely different physical environments.
Introduction.-Understanding the nature of supradense nuclear matter and its EOS, encoded in the pressure–energy density relation $P(\varepsilon)$, is a longstanding challenge shared by nuclear physics and astrophysics. In the gravitational regime, the $P(\varepsilon)$ relation determines the internal structure of neutron stars through the Tolman–Oppenheimer–Volkoff (TOV) equations [1, 2]. Recent progress has shifted attention toward constraining the dimensionless trace anomaly $\Delta\equiv 1/3-P/\varepsilon$ [3], which characterizes deviations from conformal behavior and the approach to the asymptotic limit predicted by perturbative QCD at high densities [4, 5, 6, 7, 8, 9, 10, 11, 12].
@@
In this Letter, we provide such an independent test by performing a Bayesian inference of the cold dense-matter trace anomaly from collective flow observables [14, 15, 16] in heavy-ion collisions. Collective flow encodes the integrated response of strongly interacting matter to pressure gradients over the full space–time evolution of the collision. We show that the trace anomaly extracted from laboratory data at GSI (FOPI [17] and HADES [18]) is quantitatively consistent with constraints inferred from GW170817 and NICER observations. This agreement identifies $\Delta(\varepsilon)$ as a unifying macroscopic property linking femtometer-scale nuclear dynamics to kilometer-scale neutron star structure. While trace-anomaly–related quantities have been extensively studied in high-temperature quark–gluon plasma created at ultrarelativistic energies, this work presents the first Bayesian extraction of the cold dense-matter trace anomaly from collective flow in heavy-ion collisions and its quantitative agreement with independent neutron star constraints.
Intrinsic composition degeneracy of TOV equations and neutron star global observables.- The TOV equations (with $c=1$) [1, 2]
$\frac{\mathrm{d}}{\mathrm{d}r}P=$ $-\frac{G(\varepsilon+P)(M+4\pi r^{3}P)}{r^{2}(1-2GM/r)},~~\frac{\mathrm{d}}{\mathrm{d}r}M=4\pi r^{2}\varepsilon$ (1)
@@
govern the internal structure and global observables of neutron stars. Especially since GW170817, great progress has been made not only in narrowing down the $P(\varepsilon)$ band, but also more recently, in constraining the dimensionless ratio $\phi\equiv\frac{P}{\varepsilon}$ which determines the trace anomaly $\Delta(\varepsilon)$ [3]. With sound speed squared $c_{s}^{2}(\varepsilon)\equiv\frac{dP}{d\varepsilon}$, the ratio $\phi$ represents the energy-density–averaged sound speed squared $\langle c_{s}^{2}(\varepsilon)\rangle$ [19, 20]:
$\langle c_{s}^{2}(\varepsilon)\rangle=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}c_{s}^{2}(\varepsilon^{\prime})\,\mathrm{d}\varepsilon^{\prime}=\frac{P(\varepsilon)}{\varepsilon}=\phi(\varepsilon).$ (2)
@@
Perturbative QCD (pQCD) predicted that $\Delta$ approaches zero from above at asymptotically high densities, reflecting the approximate restoration of conformal symmetry of quark matter [4, 5, 6, 7, 8, 9, 10, 11, 12]. Moreover, General Relativity (GR) sets a lower limit of $\Delta_{\textrm{G}R}\geq-(0.041\sim 0.051)$ [21, 22, 23]. These two limits are indicated in Fig. 1.
@@
Figure 1: Trace anomaly $\Delta(\varepsilon)$ as a function of reduced energy density. Values are shown as inferred from Bayesian analyses of neutron star observational data by (1) Gorda et al. [8], (2) Al-Mamun et al. [24], and (3) Fujimoto et al. [3]. Also plotted are the central $\Delta(\varepsilon)$ values [25] for the binary neutron stars GW170817 [26], PSR J0030+0451 [28, 27], and PSR J0740+6620 [29]. These astrophysical constraints are compared with Bayesian-inferred $\Delta(\varepsilon)$ values (green squares) obtained from proton collective flow data measured by the FOPI and HADES Collaborations at GSI [17, 18]. For historical context, we also include $\Delta(\varepsilon)$ values obtained by converting the continuous pressure bands previously extracted through forward-modeling of kaon production [30, 31] and collective flow at BEVALAC and AGS energies [32]. The pQCD and GR limits, as well as $\Delta(\rho_{0})=1/3$ are indicated by the horizontal lines.
@@
The TOV equations are intrinsically composition degenerate, namely, they only see the $P(\varepsilon)$ but not its microscopic origin. Whether the pressure comes from nucleons, hyperons, quarks, two-body forces, three-body forces, short-range correlations, or other sources is irrelevant unless they uniquely change the EOS itself. While in forward modeling of neutron stars, one can use a chosen microscopic model to construct a unique EOS to predict observables (e.g., mass, radius, tidal deformation), the inverse inference from observables to microscopic models is not unique. Fundamentally, the forward modeling tells us how a given microscopic theory maps into observables. But inverse inference tells us what part of the theory survives coarse-graining. That surviving part from analyzing neutron star global observables is $P(\varepsilon)$ and its derivatives, but not the underlying Hamiltonian. Thus, neutron star observables constrain the macroscopic EOS, while each forward modeling provides one possible microscopic realization consistent with it. Unlike the EOS $P(\varepsilon)$ itself, the trace anomaly $\Delta$ is dimensionless, insensitive to absolute energy scales and additive constants in $\varepsilon$, and weakly sensitive to how kinetic vs potential energy is split. Moreover, because many microscopic differences modify $P$ and $\varepsilon$ in correlated ways, these variations largely cancel in the ratio $\phi$, making $\Delta(\varepsilon)$ significantly more model-independent than the EOS itself.
@@
Indeed, the $\Delta(\varepsilon)$ inferred recently from neutron star observations using various Bayesian frameworks constrained by chiral effective field theory at low densities and perturbative QCD at high densities, yielded mutually consistent results across several independent analyses as illustrated by three examples in Fig. 1 from Gorda et al. [8], Al-Mamun et al. [24] and Fujimoto et al. [3]. Shown also are the central $\Delta(\varepsilon)$ values [25] in the binary neutron stars in GW170817 [26], PSR0030+0451 [28, 27] and PSR0740+6620 [29], inferred without using any input EOS model from perturbative solutions of the dimensionless TOV equations for reduced NS variables [21]. Thus, the neutron star data do not select a unique microscopic Hamiltonian but instead reveal a universal, model-independent trace anomaly of dense matter.
@@
What can be Bayesian inferred from nuclear collective flow data in heavy-ion collisions?— Collective flow observables [14, 15, 16] in intermediate-energy heavy-ion collisions are commonly interpreted within hydrodynamic or transport frameworks in which pressure gradients drive the acceleration of the medium [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. At the level of ideal relativistic hydrodynamics, this physics [15, 43] is encapsulated in the Euler equation governing the collective velocity field $\vec{v}$,
@@
where $\varepsilon+P$ plays the role of an inertial mass density and $\nabla P$ provides the driving force. Importantly, Eq. (3) depends only on the macroscopic equation of state $P(\varepsilon)$ and is insensitive to the microscopic mechanisms that generate the pressure and energy density. We emphasize that Eq. (3) should be understood as a macroscopic organizing principle rather than a statement that the system involved in intermediate energy heavy-ion collisions behaves as an ideal fluid.
All microscopic physics—including many-body forces, short-range correlations, momentum dependence of the single-nucleon potential, in-medium scattering cross sections, or possible changes in the relevant degrees of freedom—affects the collective flow only insofar as it modifies the functional relation between $P$ and $\varepsilon$. In direct analogy with the TOV equations for neutron stars, the Euler equation therefore guarantees that collective flow observables constrain the coarse-grained EOS rather than the specific microscopic Hamiltonians. In Bayesian analyses of collective flow, microscopic transport simulations are used to generate training data for emulators that span a sufficiently broad space of EOS realizations. The inference itself, however, is limited to macroscopic EOS properties that enter Eq. (3). Consequently, flow observables cannot uniquely determine the microscopic composition or interaction structure of dense matter, but instead constrain the coarse-grained EOS $P(\varepsilon)$ and its derivatives, which govern the macroscopic response of the system.
Moreover, collective flow is generated continuously throughout the entire space–time evolution of a heavy-ion collision, as demonstrated quantitatively in, e.g., Refs. [44, 45]. Pressure gradients acting at different times and spatial locations accelerate the medium locally, and the observed final-state flow represents an integral over the full dynamical history of the system. As a result, flow observables do not probe the EOS at a single energy density, but encode a weighted average of the medium response over the range of energy densities explored during the reaction.
This averaging property can be made explicit by rewriting Eq. (3) as
@@
and integrating over the space–time evolution of the collision. The resulting collective flow therefore reflects a space–time–weighted average of the sound speed squared $c_{s}^{2}(\varepsilon)$, leading to a stronger correlation with the trace anomaly $\Delta(\varepsilon)$ than with strictly local values of $c_{s}^{2}(\varepsilon)$ at any given time. According to Eq. (2), distinct microscopic models with substantially different local $c_{s}^{2}(\varepsilon)$ profiles can thus yield the same trace anomaly $\Delta(\varepsilon)$.
Bayesian inference of the trace anomaly and speed of sound from collective flow in heavy-ion collisions.— Using a Gaussian-process emulator trained on an isospin-dependent Boltzmann–Uehling–Uhlenbeck (IBUU) transport model [46, 47] for intermediate-energy heavy-ion reactions, we recently performed Bayesian analyses [48, 49] of the excitation functions of protons’ directed and elliptic flow measured by the FOPI Collaboration at beam energies from 150 to 1200 MeV/nucleon, as well as the HADES data at 1230 MeV/nucleon [18], for mid-central Au+Au collisions. In transport simulations of heavy-ion collisions, the EOS of cold nuclear matter enters through the single-nucleon mean-field potential, while temperature-related effects are treated separately via collision integrals [34]. To account for uncertainties associated with particle scattering in dense matter, we introduced an in-medium baryon–baryon scattering cross-section (BBSCS) modification factor $X$ (relative to free-space values) and assign it a broad prior range of $0.3\leq X\leq 2.0$.
Figure 2: The averaged speed of sound squared $c^{2}_{s}(\rho)>=P/\varepsilon=\phi$ as a function of reduced baryon density with a Skyrme-like EOS with varying incompressibility parameter $K$ (upper panel) and nuclear symmetry energy (lower panel), respectively.
The mean-field potential for a baryon $q$ is parameterized in a Skyrme-like form
$V_{q}(\rho,\delta)=a\left(\frac{\rho}{\rho_{0}}\right)+b\left(\frac{\rho}{\rho_{0}}\right)^{\sigma}+V_{\textrm{s}ym}^{q}(\rho,\delta)+V^{q}_{\textrm{C}ou},$ (5)
where $V_{\textrm{s}ym}^{q}(\rho,\delta)$ denotes the symmetry potential in isospin-asymmetric nuclear matter with isospin asymmetry $\delta=(\rho_{n}-\rho_{p})/\rho$, and $V^{q}_{\textrm{C}ou}$ is the Coulomb potential. The parameters $a$, $b$, and $\sigma$ are fixed by the saturation properties of symmetric nuclear matter (SNM) at density $\rho_{0}\simeq 0.16~\mathrm{fm}^{-3}$. By varying the single incompressibility parameter
$K=9\rho_{0}^{2}\left[\frac{d^{2}E(\rho,\delta)}{d\rho^{2}}\right]_{\rho_{0},\delta=0},$
one systematically scans the full stiffness range of the SNM EOS at suprasaturation densities.
The corresponding energy per nucleon $E(\rho,\delta)$ in isospin-asymmetric nuclear matter is given by
$E=\frac{3}{5}E_{F}^{0}\left(\frac{\rho}{\rho_{0}}\right)^{2/3}+\frac{a}{2}\frac{\rho}{\rho_{0}}+\frac{b}{1+\sigma}\left(\frac{\rho}{\rho_{0}}\right)^{\sigma}+E_{\textrm{s}ym}(\rho)\,\delta^{2},$ (6)
where $E_{F}^{0}$ is the nucleon Fermi energy at $\rho_{0}$ and $E_{\textrm{s}ym}(\rho)$ is the nuclear symmetry energy. In analyzing the GSI flow data, we adopt a conservative prior range K=180K=180–400 MeV. As shown in the upper panel of Fig. 2, for a typical isospin asymmetry $\delta\simeq 0.2$ achievable in heavy-ion collisions and a linear $E_{\textrm{s}ym}(\rho)$ motivated by relativistic mean-field models, the ratio $\phi=P/\varepsilon$, with $P(\rho)=\rho^{2}\,dE(\rho,\delta)/d\rho$ and $\varepsilon=\rho E(\rho,\delta)$, spans a wide range at suprasaturation densities. Moreover, as illustrated in the lower panel of Fig. 2, varying $E_{\textrm{s}ym}(\rho)$ within its current empirical uncertainties [50] induces only a weak effect on $\phi$, reflecting the dominance of the SNM EOS in both pressure and energy density.
Finally, it is worth emphasizing that the absence of explicit momentum dependence in the single-nucleon potential allows a clean decoupling of the cold-matter EOS from thermal effects during the collision dynamics. As discussed above, the inverse Bayesian inference of the trace anomaly from collective flow is largely insensitive to microscopic composition and interaction details, and instead constrains the macroscopic EOS regardless of how it is constructed.
Table 1: Beam energy dependence of the most probable $K$ and $X$ values inferred from the Bayesian analyses of FOPI’s excitation functions of protons’ directed and elliptic flow data [49].
$E_{\textrm{beam}}/A$ (MeV) $K$ (MeV) X
150 $220\pm 55$ $0.7^{+1.1}_{-0.4}$
250 $320\pm 62$ $0.6^{+1.3}_{-0.5}$
400 $313\pm 63$ $0.9^{+1.3}_{-0.8}$
600 $312\pm 60$ $0.9^{+1.4}_{-0.8}$
800 $309\pm 63$ $0.9^{+1.3}_{-0.7}$
1000 $310\pm 62$ $0.9^{+1.3}_{-0.7}$
1200 $323\pm 53$ $1.0^{+1.4}_{-0.9}$
Listed in Table LABEL:mean are the most probable values of the incompressibility $K$ and the in-medium scattering factor $X$ inferred from our Bayesian analyses of the beam-energy dependence of protons’ directed and elliptic flow. The corresponding time evolution of the central baryon density, $\rho_{\textrm{c}en}/\rho_{0}$, evaluated in a cubic cell of volume $1~\mathrm{fm}^{3}$ centered at the center of mass of the reaction system, is shown in Fig. 3. For each beam energy, the maximum central density attained during the collision is used to evaluate the trace anomaly, with the sound speed squared $c_{s}^{2}$ averaged up to this density according to Eq. (2). The systematic beam-energy dependence of the maximum central density then enables the extraction of the trace anomaly as a function of the reduced energy density, shown as green squares in Fig. 1.
Figure 3: Time evolution of central baryon density in mid-central Au+Au reactions with beam energies from 150 to 1200 MeV/nucleon with parameters listed in Table LABEL:mean.
Remarkably, the trace anomaly extracted from collective flow in heavy-ion collisions exhibits quantitative agreement with independent constraints inferred from neutron star observations based on mass–radius and tidal deformability measurements. This agreement arises because both systems probe the same coarse-grained EOS $P(\varepsilon)$ through integrated macroscopic responses, while remaining largely insensitive to the microscopic composition of the pressure and energy density. In the Bayesian analysis, the specific microscopic transport model used in the forward simulations serves solely as a generator of training data that spans a sufficiently broad space of EOS realizations—most notably encompassing the full range of the incompressibility parameter KK—rather than as a source of model-dependent microscopic constraints. Consequently, the inferred trace anomaly reflects EOS information that is robust under changes of microscopic dynamics, explaining its consistency with independently inferred neutron star constraints.
@@
While the matter produced in heavy-ion collisions at intermediate beam energies does not behave as a perfect ideal fluid, the transport-model framework employed here naturally incorporates non-ideal effects such as finite mean free paths and viscosity, reinforcing the robustness of the inferred trace anomaly beyond the ideal-hydrodynamic limit.
@@
For comparison, we have also translated the $P(\rho)$ bands extracted approximately two decades ago, from kaon production data analyzed within quantum molecular dynamics (QMD) transport models [30, 31], and from collective flow data at BEVALAC and AGS energies studied using both QMD- and BUU-type transport models [32], into the trace anomaly $\Delta(\varepsilon)$ by assuming $\varepsilon=M_{N}\rho$, where $M_{N}$ is the average nucleon mass. Interestingly, these earlier constraints from forward-modeling are generally consistent with both the neutron star $\Delta(\varepsilon)$ bands and the results extracted here from the FOPI and HADES flow data, providing additional support for the universal character of the dense-matter trace anomaly.
@@
Having obtained the energy-density dependence of the trace anomaly $\Delta(\varepsilon)$, Eq. (2) can be inverted to recover the corresponding profile of the sound speed squared $c_{s}^{2}(\varepsilon)$ [3],

### v2

# Trace Anomaly of Cold Dense Matter Constrained by Collective Flow

@@ Abstract
The trace anomaly of dense matter, $\Delta\equiv 1/3-P/\varepsilon$, defined through the ratio $w\equiv P/\varepsilon$ of pressure $P$ to energy density $\varepsilon$, quantifies deviations from conformal symmetry and provides a dimensionless measure of the stiffness of the equation of state (EOS) relevant for both neutron stars and heavy-ion collisions. While $\Delta(\varepsilon)$ has recently been inferred from neutron star observations, we report the first Bayesian extraction of the trace anomaly from collective flow observables in intermediate-energy heavy-ion collisions. By employing transport-model simulations that explicitly decouple the cold matter mean-field potential from thermal effects, we directly constrain the EOS of cold dense matter. Remarkably, the trace anomaly inferred from laboratory flow data agrees quantitatively, within $68\%$ credible intervals, with independent astrophysical posterior bands. This nontrivial agreement demonstrates that heavy-ion collisions and neutron star observations probe the same macroscopic properties in a mutually consistent way, establishing the dense-matter trace anomaly as a composition-insensitive macroscopic bridge observable across widely different physical environments.
Introduction.-Understanding the nature of supradense nuclear matter and its EOS, encoded in the pressure–energy density relation $P(\varepsilon)$, is a longstanding challenge shared by nuclear physics and astrophysics. In the gravitational regime, the $P(\varepsilon)$ relation determines the internal structure of neutron stars through the Tolman–Oppenheimer–Volkoff (TOV) equations [1, 2]. Recent progress has shifted attention toward constraining the dimensionless trace anomaly $\Delta\equiv 1/3-P/\varepsilon$ [3]. The latter is uniquely determined by the dimensionless EOS parameter $w\equiv P/\epsilon$. While $w$ is traditionally used in cosmology to classify the expansion dynamics of radiation, matter, and vacuum energy [4, 5], here it serves as a scale-independent measure of the medium’s intrinsic stiffness. By adopting this universal parameter, we can more clearly quantify the approach to the conformal limit ($w\to 1/3$) predicted by perturbative QCD at high densities [6, 7, 8, 9, 10, 11, 12, 13, 14, 15], and identify consistent macroscopic patterns across the vastly different energy scales of heavy-ion collisions and neutron stars.
@@
In this Letter, we provide such an independent test by performing a Bayesian inference of the cold dense-matter trace anomaly from collective flow observables [17, 18, 19] in heavy-ion collisions. Collective flow encodes the integrated response of strongly interacting matter to pressure gradients over the full space–time evolution of the collision. We show that the trace anomaly extracted from laboratory data at GSI (FOPI [20] and HADES [21]) is quantitatively consistent with constraints inferred from GW170817 and NICER observations. This agreement identifies $\Delta(\varepsilon)$ as a unifying macroscopic property linking femtometer-scale nuclear dynamics to kilometer-scale neutron star structure. While trace-anomaly–related quantities have been extensively studied in high-temperature quark–gluon plasma created in heavy-ion collisions at ultrarelativistic energies [22], QCD thermodynamics/EOS [23, 24] and hadron physics [25], as well as in cosmology regarding the cosmological vacuum energy [26], EOS of dark energy [27], quantum gravity [28] and cosmological inflation applications [29], this work presents the first Bayesian extraction of the cold dense-matter trace anomaly from collective flow in heavy-ion collisions and its quantitative agreement with independent neutron star constraints.
@@
Intrinsic composition degeneracy of TOV equations and neutron star global observables.- Especially since GW170817, great progress has been made not only in narrowing down the $P(\varepsilon)$ band, but also more recently, in constraining the trace anomaly $\Delta(\varepsilon)$ [3]. With sound speed squared $c_{s}^{2}(\varepsilon)\equiv\frac{dP}{d\varepsilon}$, the ratio $w$ represents the energy-density–averaged value $\langle c_{s}^{2}(\varepsilon)\rangle$ up to $\varepsilon$ [30, 31]:
$\langle c_{s}^{2}(\varepsilon)\rangle=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}c_{s}^{2}(\varepsilon^{\prime})\,\mathrm{d}\varepsilon^{\prime}=\frac{P(\varepsilon)}{\varepsilon}=w(\varepsilon).$ (1)
@@
Perturbative QCD (pQCD) predicted that $\Delta$ approaches zero from above at asymptotically high densities, reflecting the approximate restoration of conformal symmetry of quark matter [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Moreover, General Relativity (GR) requires $\Delta_{\textrm{G}R}\geq-(0.041\sim 0.051)$ [32, 33, 34].
@@
Figure 1: Trace anomaly $\Delta(\varepsilon)$ as a function of reduced energy density $\varepsilon/\varepsilon_{0}$ with $\varepsilon_{0}\simeq 150~\mathrm{MeV/fm}^{3}$ denotes the energy density at the saturation nucleon density $\rho_{0}\simeq 0.16~\mathrm{fm}^{-3}$ of symmetric nuclear matter. Values are shown as inferred from Bayesian analyses of neutron star observational data by (1) Gorda et al. [10], (2) Al-Mamun et al. [37], and (3) Fujimoto et al. [3]. Also plotted are the central $\Delta(\varepsilon)$ values [38] for the binary neutron stars GW170817 [39], PSR J0030+0451 [41, 40], and PSR J0740+6620 [42]. These astrophysical constraints are compared with Bayesian-inferred $\Delta(\varepsilon)$ values (green squares) obtained from proton collective flow data measured by the FOPI and HADES Collaborations at GSI [20, 21]. For historical context, we also include $\Delta(\varepsilon)$ values obtained by converting the continuous pressure bands previously extracted through forward-modeling of kaon production [43, 44] and collective flow at BEVALAC and AGS energies [45]. The pQCD limit and $\Delta(\rho_{0})=1/3$ are indicated by the horizontal lines.
Since the trace anomaly is fully determined by the EOS $P(\varepsilon)$, one may ask why extracting it provides a more robust or less model-dependent constraint than directly inferring the EOS itself. We address this important question from two complementary aspects. First, the ratio $w=P/\varepsilon$ is a dimensionless quantity that characterizes the intrinsic stiffness of the EOS. Dimensionless EOS parameters are often more robust against microscopic modeling uncertainties because they remove normalization scales associated with the absolute values of pressure and energy density. A well-known example is the use of $w$ in cosmology to classify radiation, matter, and vacuum energy independently of microscopic details [4, 5]. Similarly, in hot QCD thermodynamics the trace anomaly $(\varepsilon-3P)/T^{4}$ is widely used to quantify interaction effects in strongly interacting matter at temperature $T$ [23, 24, 25]. Analogously, the ratio $P/\varepsilon$ in dense matter provides a dimensionless, averaged measure of the EOS stiffness, as shown in Eq. (1), and therefore provides a robust characterization of the thermodynamic response of matter under extreme compression.
Second, the robustness of the ratio $P/\varepsilon$ can be understood analytically using the Hugenholtz–Van Hove (HVH) theorem [35, 36]. For a zero-temperature Fermi system, $P/\varepsilon=\rho E_{\textrm{F}}/\varepsilon-1=E_{\textrm{F}}/{E}-1$, depends only on the relation between the Fermi energy $E_{\textrm{F}}$ and the average energy per particle $E$, and is therefore insensitive to how the energy density is partitioned into kinetic and interaction contributions. This further demonstrates that the trace anomaly provides a macroscopic characterization of the EOS independent of microscopic decomposition.
@@
It is known that the TOV equations [1, 2] are intrinsically composition degenerate, namely, they only see the $P(\varepsilon)$ but not its microscopic origin. Whether the pressure comes from nucleons, hyperons, quarks, two-body forces, three-body forces, short-range correlations, or other sources is irrelevant unless they uniquely change the EOS itself. While in forward modeling of neutron stars, one can use a chosen microscopic model to construct a unique EOS to predict observables (e.g., mass, radius, tidal deformation), the inverse inference from observables to microscopic models is not unique. Fundamentally, the forward modeling tells us how a given microscopic theory maps into observables. But inverse inference tells us what part of the theory survives coarse-graining. That surviving part from analyzing neutron star global observables is $P(\varepsilon)$ and its derivatives, but not the underlying Hamiltonian. Thus, neutron star observables constrain the macroscopic EOS, while each forward modeling provides one possible microscopic realization consistent with it.
@@
Indeed, the trace anomaly $\Delta(\varepsilon)$ inferred recently from neutron star observations using various Bayesian frameworks—constrained by chiral effective field theory at low densities and perturbative QCD at high densities—yields mutually consistent results across several independent analyses. This is illustrated by three examples in Fig. 1 from Gorda et al. [10], Al-Mamun et al. [37], and Fujimoto et al. [3]. These studies employ different EOS parameterizations, priors, and statistical implementations, yet the resulting $\Delta(\varepsilon)$ bands overlap closely within their quoted $68\%$ credible regions. Also shown are the central values of $\Delta(\varepsilon)$ [38] for the binary neutron star merger GW170817 [39], PSR0030+0451 [41, 40] and PSR0740+6620 [42] observed by NICER, inferred without assuming any specific microscopic EOS model from perturbative solutions of the dimensionless TOV equations for reduced neutron star variables [32]. Taken together, these results indicate that current neutron star observations constrain a relatively robust, coarse-grained measure of the EOS stiffness, even though the detailed inference of $\Delta(\varepsilon)$ may depend on the adopted parameterizations and priors.
@@
What can be Bayesian inferred from nuclear collective flow data in heavy-ion collisions at intermediate energies?— Collective flow observables [17, 18, 19] in intermediate-energy heavy-ion collisions are commonly interpreted within hydrodynamic or transport frameworks in which pressure gradients drive the acceleration of the medium [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. At the level of ideal relativistic hydrodynamics this physics [18, 56] is encapsulated in the Euler equation governing the collective velocity field $\vec{v}$,
@@
where $\varepsilon+P$ plays the role of an inertial mass density and $\nabla P$ provides the driving force. Importantly, Eq. (2) depends only on the macroscopic EOS $P(\varepsilon)$ and is insensitive to the microscopic mechanisms that generate the pressure and energy density.
We emphasize that Eq. (2) should be understood as a macroscopic organizing principle rather than implying that the matter created in intermediate-energy heavy-ion collisions behaves as an ideal fluid. In practice, collective flow has long been analyzed using transport approaches that incorporate finite mean free paths, non-equilibrium dynamics, and microscopic scattering processes. Within these models, a variety of microscopic inputs—such as the momentum dependence of the single-particle potential, in-medium modifications of the scattering cross sections, and different realizations of the nuclear EOS, can affect the detailed magnitude of flow observables. Nevertheless, extensive forward-modeling studies over the past four decades have shown that the pressures extracted from flow data are remarkably consistent within relatively narrow bands (with respect to those extracted from neutron star properties, see, e.g., Refs. [57, 58]) across different transport implementations, from the pioneering work of Ref. [45] to recent systematic surveys [59, 60, 61]. This empirical convergence indicates that collective flow is primarily sensitive to the macroscopic relation $P(\varepsilon)$ rather than to the specific microscopic ingredients entering the transport dynamics.
Collective flow reflects an integral response to pressure gradients over the full space–time evolution of the collision  [62, 63]. Rewriting Eq. (2) as
@@
shows that the observed flow encodes a space–time–weighted average of $c_{s}^{2}(\varepsilon)$ and is therefore primarily sensitive to coarse-grained quantities such as $\Delta(\varepsilon)$. Since $P(\varepsilon)=\int_{0}^{\varepsilon}c_{s}^{2}(\varepsilon^{\prime})\,d\varepsilon^{\prime}$, the trace anomaly $\Delta(\varepsilon)$ encodes the EOS stiffness averaged up to $\varepsilon$. Evaluated at the maximum density reached in the collision, it therefore characterizes the effective stiffness probed by the reaction.
Within this framework, using a Gaussian-process emulator trained on an isospin-dependent Boltzmann–Uehling–Uhlenbeck (IBUU) transport model [64, 65] for intermediate-energy heavy-ion reactions with an EOS described in End Matter and Refs. [66, 67], we infer the trace anomaly $\Delta(\varepsilon)$. To our knowledge, this represents the first Bayesian extraction of this quantity from intermediate-energy heavy-ion collective flow data. Future studies employing different transport models using various EOS parameterizations, including possibly meta-model EOSs (see, e.g., Ref. [68]), will provide valuable cross-checks of the robustness of this inference, although such model-to-model comparisons lie beyond the scope of the present work.
@@
Bayesian inference of the trace anomaly and speed of sound from collective flow in heavy-ion collisions.— We recently performed Bayesian analyses [66, 67] of the excitation functions of protons’ directed and elliptic flow measured by the FOPI Collaboration at beam energies from 150 to 1200 MeV/nucleon, as well as the HADES data at 1230 MeV/nucleon [21], for mid-central Au+Au collisions. The matter produced in heavy-ion collisions at intermediate beam energies does not behave as a perfectly ideal fluid; the transport-model framework employed here naturally incorporates non-ideal effects such as finite mean free paths and viscosity, reinforcing the robustness of the inferred trace anomaly beyond the ideal-hydrodynamic limit.
In transport simulations of heavy-ion collisions at intermediate energies, the EOS of cold nuclear matter enters through the single-nucleon mean-field potential[47]. Temperature-related effects are treated separately via collision integrals. In particular, to account for uncertainties associated with particle scattering in dense matter, we introduced an in-medium baryon–baryon scattering cross-section (BBSCS) modification factor $X$ (relative to free-space values) and assign it a broad prior range of $0.3\leq X\leq 2.0$. Posterior distributions are obtained using the emulator-based Bayesian framework described in Refs. [66, 67]. The posterior distributions are well localized within the adopted prior ranges and are therefore primarily data-driven. Because the inference constrains the macroscopic relation $P(\varepsilon)$ rather than a specific microscopic realization, different EOS parameterizations that reproduce the same $P(\varepsilon)$ will yield essentially identical trace anomalies within the present uncertainties. Additional discussions of the role of momentum-dependent interactions and their impact on the extraction of the cold EOS are provided in the End Matter.
The time evolution of the central baryon density, $\rho_{\textrm{c}en}/\rho_{0}$, evaluated in a cubic cell of volume $1~\mathrm{fm}^{3}$ centered at the center of mass of the reaction system, is shown in Fig. 2. For each beam energy, the maximum central density attained during the collision is used to evaluate the trace anomaly, with the sound speed squared $c_{s}^{2}$ averaged up to this density according to Eq. (1). The systematic beam-energy dependence of the maximum central density then enables the extraction of the trace anomaly as a function of the reduced energy density, shown as green squares in Fig. 1.
Figure 2: Time evolution of central baryon density in mid-central Au+Au reactions with beam energies from 150 to 1200 MeV/nucleon with parameters listed in Table LABEL:mean of End Matter.
Remarkably, as shown in Fig. 1, the trace anomaly extracted from collective flow in heavy-ion collisions exhibits quantitative agreement with independent constraints inferred from neutron star observations based on mass–radius and tidal deformability measurements. This agreement provides an independent consistency check of the EOS. In the Bayesian analysis, the specific microscopic transport model used in the forward simulations serves solely as a generator of training data that spans a sufficiently broad space of EOS realizations—most notably encompassing the full range of the incompressibility parameter KK—rather than as a source of model-dependent microscopic constraints. Consequently, the inferred trace anomaly reflects EOS information that is robust under changes of microscopic dynamics, explaining its consistency with independently inferred neutron star constraints.
@@
For comparison, we have also translated the $P(\rho)$ bands extracted approximately two decades ago, from kaon production data analyzed within quantum molecular dynamics (QMD) transport models [43, 44], and from collective flow data at BEVALAC and AGS energies studied using both QMD- and BUU-type transport models [45], into the trace anomaly $\Delta(\varepsilon)$. In making this comparison, we approximate the energy density as $\varepsilon=M_{N}\rho$ where MN is the mean mass of nucleons, since the earlier constraint bands were obtained by combining multiple analyses based on different nuclear interactions. The resulting uncertainty introduces only a very small horizontal shift because the combined kinetic and potential energy contributions amount to less than $\sim 10\%$ of $M_{N}$ (for $K=300$ MeV) up to $3\rho_{0}$. We emphasize that in our Bayesian analyses the energy density $\varepsilon$ includes the nucleon rest-mass, kinetic, and potential energy contributions exactly as in Eq. (7) given in End Matter. These earlier constraints are broadly consistent with both the neutron star $\Delta(\varepsilon)$ bands and the results extracted here from the FOPI and HADES flow data, supporting the robustness of the trace anomaly as a measure of dense-matter stiffness.
@@
Because $P(\varepsilon)=\int_{0}^{\varepsilon}c_{s}^{2}(\varepsilon^{\prime})\,d\varepsilon^{\prime}$, one may interpret $P/\varepsilon$ as an effective average of $c_{s}^{2}$ over density based on the mean value theorem. This non-uniqueness does not affect the physical interpretation. By studying the beam-energy dependence of the flow signal, which corresponds to different maximum compression densities, and inverting Eq. (1), we reconstruct the density dependence of the sound speed squared $c_{s}^{2}(\varepsilon)$. Analytically, the density dependence of the sound speed can be written directly in terms of the trace anomaly. Using the definition $\Delta(\varepsilon)=1/3-P(\varepsilon)/\varepsilon$, one obtains [3]
@@
where $\bar{\varepsilon}\equiv\varepsilon/\varepsilon_{0}$. The first term represents the derivative contribution associated with the density variation of the trace anomaly, while the remaining terms form the non-derivative part.